OpenStudy (anonymous):

Change to a percent: 4 3/8 Round to the nearest tenth of one percent.

OpenStudy (ash2326):

We have \[4\frac 38\] We need to convert into a percent \[4\frac 38=4+\frac 38=\frac {4\times8}{8}+\frac 38=\frac{32+3}{8}=\frac{35}{8}\] Now conversion to percent \[\frac{35}{8}\times 100=\] Can you do now, it's just calculation

OpenStudy (ash2326):

@elizabeth1959